Solve by the substitution method.
4x+4y=24
-5x-y=24
The solution type in ordered pair.

Answer by mananth(16949)

4x+4y=24 .............1
y=-24-5x ---------2
Plug the value of y in (1)
4x+4(-24-5x)=24
4x-96-20x=24
4x-20x=24+96
-16x=120
/-16
x=-7.5
Plug the value of x in (1)
4x+4y=24
4*-7.5 +4y=24
-30+4y=24
4y=54
/4
y=13.5